In order to find a 86% confidence interval we need to find values a and b such that for Z∼N(μ=0,σ=1),

P(a<Z<b)=0.86.

(a) Suppose a=−2.2475. Then b=
Answer #1

Z∼N(μ = 0,σ = 1),

P(a < Z < b) = 0.86.

a = - 2.2475

We have to find value of b.

P(a < Z < b) = 0.86

P( - 2.2475 < Z < b) = 0.86

P(Z < b) - P(Z < - 2.2475) = 0.86

P(Z < b) - 0.0123 = 0.86

P(Z < b) = 0.86 + 0.0123

P(Z < b) = 0.8723

b = 1.1374